# Neural Fields with Thermal Activations for Arbitrary-Scale Super-Resolution

### 29 Nov 2023 v1

@@ Abstract
Recent approaches for arbitrary-scale single image super-resolution (ASSR) have used local neural fields to represent continuous signals that can be sampled at different rates. However, in such formulation, the point-wise query of field values does not naturally match the point spread function (PSF) of a given pixel. In this work we present a novel way to design neural fields such that points can be queried with a Gaussian PSF, which serves as anti-aliasing when moving across resolutions for ASSR. We achieve this using a novel activation function derived from Fourier theory and the heat equation. This comes at no additional cost: querying a point with a Gaussian PSF in our framework does not affect computational cost, unlike filtering in the image domain. Coupled with a hypernetwork, our method not only provides theoretically guaranteed anti-aliasing, but also sets a new bar for ASSR while also being more parameter-efficient than previous methods.
@@ 1 Introduction
For years, learning-based image super-resolution (SR) methods have achieved ever better results. But unlike interpolation methods, which allow us to resample a given image at any resolution, such methods require retraining for each desired scaling factor. In recent years, several supervised methods for arbitrary-scale SR (ASSR) have been proposed to allow end-users to choose any scaling factor, rendering such systems much more flexible [17]. Notably, Chen et al. [11] pioneered the use of neural fields for single-image super-resolution (SR) with LIIF, exploiting the continuous representation to enable SR at arbitrary scaling factors. Their method has inspired several follow-ups which build upon the idea of using local neural fields [22, 8, 10]. This is not surprising: Neural fields are in many ways a natural match for variable-resolution computer vision and graphics [43]. By implicitly parameterizing a target signal as a neural network that maps coordinates to signal strength, they offer a compact representation, defined over a continuous input domain, and analytically differentiable. Among others, they have been widely adopted for 3D reconstruction with radiance [30], occupancy [29, 35], and distance fields [33, 37].
@@
Figure 1: Given a low-resolution input, a hypernetwork predicts the parameters of a customized neural field $\Phi_{\bm{\theta}}(x,y,t)$ composed of sinusoidal components. We obtain Gaussian-filtered (anti-aliased) versions of the image that correspond to arbitrary pixel sizes by simply modulating the components according to their frequencies. This representation can then be queried at arbitrary coordinates $(x,y,t)$ to recover pixel grids at any desired scaling factor.
@@
Neural fields offer a compelling representation of continuous signals. However, for many tasks it is not desirable to query such fields at a single point, but rather to have an integral observation model such as a point spread function (PSF) [3]. This is not trivial [3, 4, 16, 5]. For ASSR, where the signal should be observed at various scales and is prone to aliasing, a neighborhood-dependent observation model is desirable. Using appropriate observation modeling has not only the potential to reduce aliasing but also to more accurately reflect the physical processes of image acquisition. For this reason, Chen et al. [11] already took a first step towards learning multiscale representations through cell encoding. For radiance fields, Mip-NeRF [3] and subsequent work [4, 16, 5] have explored ways to mitigate aliasing effects by focusing on the positional encoding of the location $\mathbf{x}$ as a tool to control different frequencies in $\Phi$.
@@
In our work, we combine recent advances in implicit neural representations with ideas from classical signal theory to develop a neural field that guarantees anti-aliasing by construction. Furthermore, this filtering incurs no additional overhead: The cost of querying a field’s value does not depend on the size of the anti-aliasing filter. The key insight is that when using sinusoidal activation functions [38], one can decay different components at different rates based on their frequencies, according to Fourier theory. Inspired by the connection between Fourier decompositions and the heat equation, we propose Thera, an ASSR method with neural fields that uses a novel thermal activation function (see Fig. 1). This formulation makes it possible to analytically compute Gaussian-blurred versions of a field $\Phi(\mathbf{x})$ for any desired (isotropic) blur radius. This is equivalent to querying $\Phi$ with a Gaussian PSF. Thera outperforms competing methods in most of our experiments, is far more parameter-efficient, and provides theoretical guarantees w.r.t. anti-aliasing and multi-scale representation. We demonstrate that a lightweight variant of our algorithm still outperforms most previous methods while using 2-4 orders of magnitude fewer parameters. To provide theoretical guarantees for our thermal activations, certain constraints are imposed on the field’s neural architecture. Fortunately, these constraints can easily be overcome with the help of a hypernetwork [14], so that our method can seamlessly integrate a wide range of well-proven SR backbones. In summary, our main contributions are:
@@
    We introduce thermal activations, which enable neural fields with a principled Gaussian observation model, and as a consequence anti-aliasing with minimal overhead.
@@
    We use thermal activations to build Thera, a novel method for ASSR which is superior to previous methods while being more parameter efficient and providing theoretical guarantees w.r.t. multiscale representation.
@@ 2 Related work
Arbitrary-scale SR. In recent years, research on arbitrary-scale super-resolution (ASSR) methods has gained attention. In ASSR the scaling factor can be chosen at inference time to be any positive number, which allows for maximum flexibility, such as that of interpolation methods. The first work along this line is MetaSR [17], which infers the parameters of a convolutional upsampling layer using a hypernetwork, conditioned on the desired scaling factor. An influential successor work is LIIF [11], in which the high-resolution image is implicitly described by a neural field that is locally conditioned with features extracted from the low-resolution image. The continuous nature of the neural field allows for sampling target pixels at arbitrary locations, and thus also arbitrary resolution. Following work has since built upon the LIIF framework. For instance, UltraSR [45] improves modeling of high-frequency textures with periodic positional encodings of the coordinate space, as is common practice for neural radiance fields. LTE [22] makes learning higher frequencies more explicit by effectively implementing a learnable coordinate transformation into 2D Fourier space, prior to a forward pass through an MLP. Vasconcelos et al. [41] use neural fields in CUF to parameterize continuous upsampling filters, which enables arbitrary-scale upsampling. Most recently, methods have integrated attention mechanisms into their implicit representation, enabling the neural field to adaptively attend to image features depending on the query coordinates and scale factor, improving reconstruction quality [8, 10]. These methods provide the required scaling factor explicitly as an input to the field, thus attempting to learn the observation model from data. In this paper, we explore a way to directly integrate such a model into the neural field representation.
@@
Implicit neural representations. A neural field is a neural network trained to map real-valued coordinates onto a field’s values. Recently, neural fields have been used for parameterizing various types of visual data, including images [20, 38, 39, 11, 22, 12], 3D scenes (e.g., represented as signed distance fields [33, 38, 37, 42], occupancy fields [29, 35], LiDAR fields [19], view-dependent radiance fields [30, 3, 4, 5]), or digital humans [46, 49, 9, 44, 13]. We refer the interested reader to Xie et al. [43] for a comprehensive survey of neural fields in vision and beyond. Frequently, it is desirable to impose some prior over the space of learnable representations. A common approach for such conditioning of neural fields is encoder-based inference [43], where a parametric encoder maps input observations to a set of latent codes $\bm{z}$, which are often local [11, 22, 41, 8, 10]. The encoded latent variables $\bm{z}$ are then used to condition field parameters $\bm{\theta}$, for instance by simply concatenating $\bm{z}$ to the coordinate inputs or through a more expressive hypernetwork  [14]. The latter is a trainable neural network $\Psi\colon\bm{z}\mapsto\bm{\theta}$, mapping latent codes to field parameters. An early example of this is Sitzmann et al. [38], and this approach is gaining popularity [43].
@@
Anti-aliasing in neural fields. Early in the recent development of implicit neural representations, concerns regarding aliasing were raised. Barron et al. [3] proposed integrating the positional encoding with Gaussian weights, which reduced aliasing in NeRF. Improvements were later proposed for unbounded scenes [4] and to improve efficiency [16]. Barron et al. [5] tackle anti-aliasing for the increasingly popular Instant-NGP [31] approach. In ASSR, methods usually provide the desired scale as input and expect to learn it from data [11, 8].
@@ 3 Method
To understand how anti-aliasing can be computed analytically, we must first define a sampling model and recap a well-established result from Fourier analysis. While we stick to the 2D case here for simplicity, all theory is valid for signals with any number of dimensions. We refer the reader to [32] for an in-depth analysis of sampling and aliasing.
To minimize aliasing when sampling a continuous signal $f(\mathbf{x})$ in 2D image space, $\mathbf{x}\in\mathbb{R}^{2}$, it is necessary to suppress high-frequency parts of the signal that cannot be represented at the desired sampling rate. One must apply a low-pass filter $g(\mathbf{x})$ whose cut-off frequency is aligned with the Nyquist frequency of the sampling rate, and then sample the band-limited signal $f\circledast g(\mathbf{x})$. Here, we show that one can naturally integrate this behavior into a neural field’s architecture. The key to this is knowing that if a signal is decomposed into sinusoidal components, filtering can be done simply by re-scaling each component by a frequency-dependent factor. This equivalence is depicted in Fig. 2. While decomposing a given signal into sinusoidal components is normally a costly operation, we show that we can avoid it by designing a neural field in which the decomposition itself is the native representation of the signal.
@@ 3.1 Thermal activation functions
The intuition behind our method goes back to the starting point of Fourier analysis: the heat equation. Other implicit representation methods, such as SIREN [38], optimize the parameters $\bm{\theta}_{\Phi}$ of a field $\Phi(\mathbf{x})$ by minimizing a data fidelity term $\mathbb{E}_{(\mathbf{x}_{i},\mathbf{y}_{i}\sim\mathcal{D})}\left[\mathcal{L}\left(\Phi(\mathbf{x}_{i}),\mathbf{y}_{i}\right)\right]$ over a dataset $\mathcal{D}$. Our key insight is the following: What if we take this field to represent the temperature distribution on a surface at a given time $t_{0}$? It is well known how to solve the heat equation forwards in time – its solution (on an isotropic surface) is given by a Gaussian blur of the state at $t_{0}$, where the blur radius depends on the time difference $t-t_{0}$. Our field $\Phi(\mathbf{x},t)$ now has an additional temporal dimension, but given a single slice $\Phi(\mathbf{x},t_{0})$ there is a unique solution $\Phi(\mathbf{x},t)$ for every $t\geq t_{0}$ such that the field obeys the heat equation
@@
where $\kappa>0$ is the thermal diffusivity of the medium. The solution for $0\leq t<t_{0}$ is, however, ill-posed, mirroring the fact that image downsampling is well-posed, but super-resolution is not.
@@
A naive attempt to ensure that $\Phi$ respects the heat equation in Eq. 1 would be to include an additional loss term: $\left|\frac{\partial\Phi}{\partial t}-\kappa\cdot\nabla_{\mathbf{x}}^{2}\Phi\right|$. Despite our best efforts, this approach was unsuccessful. Nevertheless, it turns out that we can design the field’s architecture in such a way that it is guaranteed to respect Eq. 1. We do so by formulating $\Phi$ as a two-layer perceptron
$\mathbf{y}=\mathbf{W}_{2}\cdot\xi(\mathbf{W}_{1}\mathbf{x}+\mathbf{b}_{1},\nu(\mathbf{W}_{1}),\kappa,t)+\textbf{b}_{2},$ (2)
where $\nu$ refers to the wave numbers (frequencies) implied by $\mathbf{W}_{1}$, described in more detail later; and the thermal activation function $\xi(\cdot)$ models the decay of different sinusoidal components over time. We set $t=0$ to be the (unobservable) initial thermal state of a given system represented as a sum of sinusoidal components:
@@
To obtain $\left|\nu(\mathbf{W}_{1})\right|$ we can simply compute the row-wise Euclidean norm of $\mathbf{W}_{1}$, corresponding to the magnitudes of the implied wave numbers (direction is irrelevant, as all components with the same frequency decay equally over time, regardless of their orientation). Equation 4 results in higher-frequency components decaying faster than lower-frequency ones, giving rise to a Gaussian low-pass filter whose strength increases with time $t$.
@@
The derivations of these values can be found in Appendix A.
@@
Refer to caption Refer to caption Refer to caption
Original Anti-aliasing for $\times$1/4 Anti-aliasing for $\times$1/16
Thera
Refer to caption Refer to caption Refer to caption
$\Phi(\mathbf{x},1)$ $\Phi(\mathbf{x},4^{2})$ $\Phi(\mathbf{x},16^{2})$
Figure 3: Comparison between a theoretical anti-aliasing filter and Thera’s anti-aliasing, which is computed without convolutions or over-sampling. Note that the field was only supervised at $\Phi(\mathbf{x},1)$.
Figure 3 shows an example where we fit the signal at $t=1$ to the image. After training, any low-pass filtered version of the image can be generated by setting $t$ according to Eq. 6. We emphasize that: (i) Computing these filtered images requires no over-sampling or convolutions; (ii) The computational cost does not depend on the size of the blur kernel or on $t$; (iii) Given $\Phi(\mathbf{x},t_{0})$, $\Phi(\mathbf{x},t)$ is known for any $t\geq t_{0}$.
By definition, such field $\Phi$ follows the heat equation and is therefore able to perform anti-aliasing at arbitrary sampling frequencies. An alternative (equivalent) interpretation is to view $\Phi(\mathbf{x},t)$ as an observation of the true, continuous signal $\Phi(\mathbf{x},0)$ with a Gaussian point spread function whose width (standard deviation) is determined by $t$.

### 14 Mar 2024 v2

@@ Abstract
Recent approaches for arbitrary-scale single image super-resolution (ASSR) have used local neural fields to represent continuous signals that can be sampled at arbitrary rates. However, the point-wise query of the neural field does not naturally match the point spread function (PSF) of a given pixel, which may cause aliasing in the super-resolved image. We present a novel way to design neural fields such that points can be queried with an adaptive Gaussian PSF, so as to guarantee correct anti-aliasing at any desired output resolution. We achieve this with a novel activation function derived from Fourier theory. Querying points with a Gaussian PSF, compliant with sampling theory, does not incur any additional computational cost in our framework, unlike filtering in the image domain. With its theoretically guaranteed anti-aliasing, our method sets a new state of the art for ASSR, while being more parameter-efficient than previous methods. Notably, even a minimal version of our model still outperforms previous methods in most cases, while adding 2-4 orders of magnitude fewer parameters. Code and pretrained models are available at https://github.com/prs-eth/thera.
@@ 1 Introduction
For years, learning-based image super-resolution (SR) methods have achieved ever better results. But unlike interpolation methods, which allow us to resample a given image at any resolution, those methods require retraining for each desired scaling factor. Recently, several schemes for arbitrary-scale SR (ASSR) have been proposed that allow end-users to choose any scaling factor, rendering the method much more flexible [17]. Notably, with LIIF, Chen et al. [11] pioneered the use of neural fields for single-image SR, exploiting their continuous representation to enable SR at arbitrary scaling factors. Their method has inspired several follow-ups which build upon the idea of using local neural fields [22, 8, 10]. This is not surprising: Neural fields are in many ways a natural match for variable-resolution computer vision and graphics [46]. By implicitly parameterizing a target signal as a neural network that maps coordinates to signal strength, they offer a compact representation, defined over a continuous input domain, and are analytically differentiable. Among others, they have been widely adopted for 3D reconstruction with radiance [32], occupancy [31, 37], and distance fields [35, 39].
@@
Figure 1: Given a low-resolution input, a hypernetwork predicts the parameters of a customized neural field $\Phi_{\bm{\theta}}(x,y,t)$ composed of sinusoidal components. We obtain Gaussian-filtered (anti-aliased) versions of the image that correspond to arbitrary pixel sizes by simply modulating the components according to their frequencies (insets illustrate components decaying at different rates). This representation can then be queried at arbitrary coordinates $(x,y,t)$ to recover pixel grids at any desired scaling factor.
@@
Neural fields offer a compelling representation of continuous signals. However, for many tasks, it is not desirable to query such fields at a single point, but rather to have an integral observation model such as a point spread function (PSF) [3]. ASSR calls for a neighborhood-dependent observation model, as the signal should be observed at various sampling frequencies and is prone to aliasing. Incorporating a physically plausible observation model is not trivial [3, 4, 25, 49, 16, 5], but has the potential to avoid aliasing. For this reason, Chen et al. [11] already took a first step towards learning multiscale representations, via cell encoding. For radiance fields, Mip-NeRF [3] and subsequent work [4, 16, 5] have explored ways to mitigate aliasing effects by focusing on the positional encoding of the input coordinates, as a means to control different frequencies in the field. Lindell et al. [25] and Yang et al. [49] have succeeded in performing subband decomposition of neural fields, though bands are architecturally fixed to a small number of discrete intervals and cannot be modified after training.
@@
In our work, we combine recent advances in implicit neural representations with ideas from classical signal theory to develop a neural field that guarantees anti-aliasing by construction. Interestingly, this filtering incurs no additional overhead: The cost of querying field values is the same for any width of the anti-aliasing filter kernel, including infinite and zero widths. The key insight is that sinusoidal activation functions [40] offer the possibility to selectively decay each component according to its spatial frequency, following Fourier theory. Inspired by the connection between Fourier decompositions and the heat equation, we propose Thera, a neural field method for ASSR underpinned by a novel thermal activation function, see Fig. 1. That formulation makes it possible to analytically compute Gaussian-blurred versions of the neural field $\Phi(\mathbf{x})$ for any desired (isotropic) blur radius. I.e., one can query $\Phi$ with a Gaussian PSF that is matched to the output resolution to prevent aliasing. [^1] While our framework is in principle not bound to the heat equation or Gaussian PSFs, it is an intuitive interpretation for readers familiar with signal theory. Thera provides theoretical guarantees w.r.t. anti-aliasing and multi-scale representation, in exchange for certain constraints on the neural architecture that can easily be overcome with a hypernetwork [14]. Empirically, it outperforms competing methods in most of our experiments and is far more parameter-efficient. We demonstrate that a lightweight variant of our model still outperforms most previous methods while adding 2-4 orders of magnitude fewer parameters. To the best of our knowledge, Thera is the first neural field method to allow bandwidth control at test time. Our main contributions are:
@@
    We introduce thermal activations, which enable neural fields with a principled Gaussian observation model, and therefore anti-aliasing with minimal overhead.
@@
    We use thermal activations to build Thera, a novel method for ASSR that offers theoretically guaranteed multi-scale capabilities, delivers state-of-the-art performance and is more parameter efficient than prior art.
@@ 2 Related Work
In ASSR the scaling factor can be chosen at inference time to be any positive number, which allows for maximum flexibility, such as that of interpolation methods. The first work along this line is MetaSR [17], which infers the parameters of a convolutional upsampling layer using a hypernetwork [14] conditioned on the desired scaling factor. An influential successor work is LIIF [11], in which the high-resolution image is implicitly described by local neural fields that are conditioned with features extracted from the low-resolution input image. The continuous nature of the neural field allows for sampling target pixels at arbitrary locations, and thus also arbitrary resolution. Following work has since built upon the LIIF framework. For instance, UltraSR [48] improves modeling of high-frequency textures with periodic positional encodings of the coordinate space, as is common practice for neural radiance fields [32, 3, 4, 5]. LTE [22] makes learning higher frequencies more explicit by effectively implementing a learnable coordinate transformation into 2D Fourier space, prior to a forward pass through an MLP. Vasconcelos et al. [43] use neural fields in CUF to parameterize continuous upsampling filters, which enables arbitrary-scale upsampling. Most recently, methods have integrated attention mechanisms into their implicit representation, enabling the neural field to adaptively attend to image features depending on the query coordinates and scale factor, improving reconstruction quality [8, 10]. These methods provide the required scaling factor explicitly as an input to the field, thus attempting to learn the observation model from data. In this paper, we explore a way to directly integrate a (Gaussian) observation model into the neural field representation.
@@
Implicit neural representations. A neural field is a neural network trained to map coordinates onto values of some physical quantity. Recently, neural fields have been used for parameterizing various types of visual data, including images [20, 40, 41, 11, 22, 12, 45], 3D scenes (e.g., represented as signed distance fields [35, 40, 39, 44, 45], occupancy fields [31, 37], LiDAR fields [19], view-dependent radiance fields [32, 3, 4, 5, 45]), or digital humans [50, 53, 9, 47, 13]. We refer the interested reader to Xie et al. [46] for a comprehensive survey of neural fields in vision and beyond. Frequently, it is desirable to impose some prior over the space of learnable representations. A common approach for such conditioning is encoder-based inference [46], where a parametric encoder maps input observations to a set of latent codes $\bm{z}$, which are often local [11, 22, 43, 8, 10]. The encoded latent variables $\bm{z}$ are then used to condition the neural field, for instance by concatenating $\bm{z}$ to the coordinate inputs or through a more expressive hypernetwork [14]. The latter is a trainable neural network $\Psi\colon\bm{z}\mapsto\bm{\theta}$, directly mapping latent codes to parameters of the field $\bm{\theta}$. An early example of this approach, which is gaining popularity [46], was proposed by Sitzmann et al. [40].
@@
Anti-aliasing in neural fields. Early in the recent development of implicit neural representations, concerns regarding aliasing were raised. Barron et al. [3] proposed integrating the positional encoding with Gaussian weights, which reduced aliasing in NeRF. Improvements were later proposed for unbounded scenes [4] and to improve efficiency [16]. Barron et al. [5] tackle anti-aliasing for the increasingly popular Instant-NGP [33] approach. Recent work has succeeded in limiting the bandwidth using multiplicative filter networks [25] or polynomial neural fields [49], although these works are restricted to discrete, pre-defined band limits (and thus resolutions) and have not tackled super-resolution tasks. Such representations would not be a good fit for ASSR since they do not allow for continuous anti-aliasing, nor bandwidth control at test time. In ASSR, methods usually provide the desired scale as input and expect to learn it from data [11, 8].
@@ 3 Method
To expound how anti-aliasing can be computed analytically, we must first define a sampling model and recap a well-established result from Fourier analysis. For simplicity, we stick to 2-dimensional signals $\mathbf{x}\in\mathbb{R}^{2}$, but all theory is valid for any number of dimensions. We refer the reader to [34] for a comprehensive discussion of sampling and aliasing.
To minimize aliasing when sampling a continuous signal $f(\mathbf{x})$ it is necessary to suppress high-frequency parts of the signal that cannot be represented at the desired sampling rate. One must apply a low-pass filter $g(\mathbf{x})$ whose cut-off frequency is aligned with the Nyquist frequency of the sampling rate, then sample the band-limited signal $f\circledast g(\mathbf{x})$. We will show that one can naturally integrate this behavior into a neural field’s architecture. The key to this is that, if a signal is decomposed into sinusoidal components, filtering can be done simply by re-scaling each component by a frequency-dependent factor. This equivalence is depicted in Fig. 2. Decomposing a given signal into sinusoidal components is normally a costly operation, but it turns out that we can avoid it by designing the neural field such that the decomposition itself is the native representation of the signal.
@@ 3.1 Thermal Activation Functions
The intuition behind our method goes back to the starting point of Fourier analysis: the heat equation. Other implicit representation methods, such as SIREN [40], optimize the parameters $\bm{\theta}$ of a field $\Phi(\mathbf{x})$ by minimizing a data fidelity term $\mathbb{E}_{(\mathbf{x}_{i},\mathbf{y}_{i}\sim\mathcal{D})}\left[\mathcal{L}\left(\Phi(\mathbf{x}_{i}),\mathbf{y}_{i}\right)\right]$ over a dataset $\mathcal{D}$. Our key insight is that we can interpret this field as the temperature distribution on an (isotropic) surface at a given time $t_{0}$. It is well known how to solve the heat equation forwards in time – its solution is obtained by blurring the initial state at $t_{0}$ with a Gaussian kernel whose radius depends on the time difference $t-t_{0}$. Our field $\Phi(\mathbf{x},t)$ now has an additional temporal dimension, but given a single slice $\Phi(\mathbf{x},t_{0})$ there is a unique solution $\Phi(\mathbf{x},t)$ for every $t\geq t_{0}$ such that the field obeys the heat equation
@@
with $\kappa>0$ the thermal diffusivity of the medium. In contrast, the solution for $0\leq t<t_{0}$ is ill-posed, mirroring the fact that image downsampling is well-posed, but super-resolution is not.
@@
It turns out that we can design the field’s architecture in such a way that it is guaranteed to respect Eq. 1. We do so by formulating $\Phi$ as a two-layer perceptron
$\Phi(\mathbf{x},t)$ $=\mathbf{W}_{2}\cdot\xi\left(\mathbf{W}_{1}\mathbf{x}+\mathbf{b}_{1},\nu(\mathbf{W}_{1}),\kappa,t\right)+\textbf{b}_{2},$ (2)
where $\bm{\theta}:=\{\mathbf{W}_{1},\mathbf{W}_{2},\mathbf{b}_{1},\mathbf{b}_{2}\}$ are MLP parameters, $\nu$ refers to the wave numbers (frequencies) implied by $\mathbf{W}_{1}$ (described in more detail later), and the thermal activation function $\xi(\cdot)$ models the decay of different sinusoidal components over time. We set $t=0$ to be the initial thermal state of a given system represented as a sum of sinusoidal components:
@@
To obtain $\left|\nu(\mathbf{W}_{1})\right|$ we can simply compute the row-wise Euclidean norm of $\mathbf{W}_{1}$, corresponding to the magnitudes of the implied wave numbers (direction is irrelevant, as all components with the same frequency decay equally, regardless of their orientation). Equation 4 results in higher-frequency components decaying faster than lower-frequency ones, giving rise to a Gaussian low-pass filter whose strength increases with time $t$.
@@
The derivations of these values can be found in Appendix 0.A.
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Original $\times$1/4 AA $\times$1/8 AA $\times$1/16 AA $\times$1/32 AA
Neural field
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\Phi(\mathbf{x},1)$ $\Phi(\mathbf{x},4^{2})$ $\Phi(\mathbf{x},8^{2})$ $\Phi(\mathbf{x},16^{2})$ $\Phi(\mathbf{x},32^{2})$
Figure 3: Comparison between a theoretical anti-aliasing filter (top row) and anti-aliasing with thermal activations (bottom row), which is computed without convolutions or over-sampling. The neural field was supervised only at $\Phi(\mathbf{x},1)$.
Figure 3 shows an example where we fit the field at $t=1$ to the image. After training, any low-pass filtered version of the image can be generated by setting $t$ according to Eq. 6. We emphasize that: (i) Computing these filtered images requires no over-sampling or convolutions; (ii) The computational cost does not depend on the size of the blur kernel or on $t$; (iii) Given $\Phi(\mathbf{x},t_{0})$, the filtered versions $\Phi(\mathbf{x},t)$ are known for any $t\geq t_{0}$.
By definition, our field $\Phi(\mathbf{x},t)$ follows the heat equation and is therefore able to perform anti-aliasing at arbitrary sampling frequencies. Note that modulating deeper fields, as is done in [30], would no longer guarantee anti-aliasing. An alternative (equivalent) interpretation is to view $\Phi(\mathbf{x},t)$ as an observation of the true, continuous signal $\Phi(\mathbf{x},0)$ with a Gaussian point spread function whose width (standard deviation) is determined by $t$.